# Certified Branch-and-Bound MaxSAT Solving (Extended Version)

### v1

@@ 22    $F\leftarrow F\cup\{\texttt{analyze}(C_{\mathcal{Q}}\vee\overline{\ell})\}$ ;
$q_{1}=\langle 3,\{y_{1}\},\{\overline{x_{1}},\overline{x_{2}}\}\rangle\text{,\qquad and}$ $q_{2}=\langle 5,\{y_{2}\},\{\overline{x_{3}},\overline{x_{4}}\}\rangle\text{.}$
@@ 0.4 Certifying CNF Encodings based on (Multi-Valued) Decision Diagrams
When a solution is found (at line 1 of Algorithm 1), MaxCDCL decides heuristically whether or not to add a CNF-encoding of the solution-improving constraint $\mathcal{O}\leq v^{*}-1$. For many CNF encodings of pseudo-Boolean constraints, it is well-known how to get certification Vandesande (2023); Gocht et al. (2022); Berg et al. (2024a). One notable exception is the CNF encoding based on binary decision diagrams (BDD), which Bofill et al. (2020) recently generalized to so-called multi-valued decision diagrams (MDD) and is used by MaxCDCL. Most of the interesting questions from a proof logging perspective already show up for BDDs, so to keep the presentation more accessible, we focus on that case and afterwards briefly discuss how this generalizes to MDDs. Formal details and proofs are included in the supplementary material.
@@
The standard way to create a reduced and ordered BDD for $\sum_{i=1}^{n}v_{i}b_{i}\leq v^{*}-1$ is to first recursively descend and create BDDs $\eta_{\mathbf{t}}$ for $\sum_{i>1}^{n}v_{i}b_{i}\leq v^{*}-a_{1}$ and $\eta_{\mathbf{f}}$ for $\sum_{i>1}^{n}v_{i}b_{i}\leq v^{*}$ and then combine them into the node $\mathrm{bdd}(x_{1},\eta_{\mathbf{t}},\eta_{\mathbf{f}})$. Using memoization, the BDD is always kept reduced: if a node is being created with the same label, and the same two children as an existing node, the existing node is returned instead; if $\eta_{\mathbf{t}}=\eta_{\mathbf{f}}$, then $\eta_{\mathbf{t}}$ is returned.
@@
$v_{\eta}$ $\Rightarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq l,$  and $v_{\eta}\Leftarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq u.$
@@
$v_{\eta}\Leftrightarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq l\text{,\quad and}$ $v_{\eta}^{\prime}\Leftrightarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq u.$
@@ 0.5 Experiments
We implemented proof logging in the MaxCDCL solver as described by Coll et al. (2025b). [^4] On top of the basic algorithm described above, the experiments include proof logging for a range of techniques included in MaxCDCL Coll et al. (2025b), namely clause vivification, equivalent and failed literal detection, clause subsumption and simplification due to unit propagation. Each solver call was assigned a single core on a 2x32-core AMD EPYC 9384X. The time and memory limits are enforced by runlim. [^5] https://github.com/arminbiere/runlim As benchmarks, the weighted instances of the MaxSAT Evaluation 2024 Berg et al. (2024b) are used. The implementation, together with the raw data from the experiments will be made available in the near future.
@@
A scatter plot evaluating the overhead in the solving time due to proof logging is shown in Figure 1. The experiments ran with a time limit of 1 hour and a memory limit of 32GB. Out of the 353 instances that were solved without proof logging, three instances could not be solved with proof logging. Considering instances solved both with and without proof logging, the median overhead of proof logging is $23\%$, which shows that for most instances, proof logging overhead is manageable. However, we see that for $10\%$ of the instances, the overhead of proof logging is larger than $153\%$. This is in line with earlier work on certifying PB-to-CNF encodings using VeriPB, such as the work of Berg et al. (2024a). Preliminary tests indicate as possible explanation for this phenomenon that writing the defining constraints for a node in an MDD is linear in the number of objective literals. In practice, this leads to large overhead for instances with a large number of objective literals. Possible ways to circumvent this is part of planned future work.
@@
While optimizing the checking time for the produced proofs was out of scope for this work, we also evaluated it by running the VeriPB proof checker VeriPB with a time limit of 10 hours and a memory limit of 64GB on the produced proofs. Out of 350 instances produced proofs, 237 were checked successfully; for the other 113 instances, the proof checker ran out of time. In the median proof verification took $23.25$ times the solving time. We believe this checking overhead to be too high for making this usable in practice. However, there are several ways in which this can be improved. On the one hand, a new proof checker is currently under development for improved performance PBOxide (2025). Simultaneously, there are ongoing efforts for building a proof trimmer for VeriPB proofs. The proofs generated by our tools can serve as benchmarks to further guide this ecosystem of VeriPB tools. On the other hand, once the proof checker is in a stable state, we can investigate which potential optimizations to the generated proofs benefit the proof checking time. Some things that come to mind are (1) using RUP with hints instead of plain RUP so that the checker does not need to perform unit propagation, and (2) deleting intermediate constraints used in the the MDD encoding to optimize memory usage of the checker. [^6] For more information on these rules, see the VeriPB repository VeriPB .
@@ 0.6 Conclusion
As future work, we plan to add proof logging to literal unlocking, a recent technique by Li et al. (2025) that improves lower bounds during look-ahead.

### v2

@@ 22    $F\leftarrow F\cup\{\texttt{analyze}(C_{\mathcal{Q}}\vee\overline{\ell})\}$ ;
$q_{1}$ $=\langle 3,\{y_{1}\},\{\overline{x_{1}},\overline{x_{2}}\}\rangle\text{ \qquad and}$
$q_{2}$ $=\langle 5,\{y_{2}\},\{\overline{x_{3}},\overline{x_{4}}\}\rangle\text{.}$
@@ 0.4 Certifying CNF Encodings based on (Multi-Valued) Decision Diagrams
When a solution is found (at line 1 of Algorithm 1), MaxCDCL decides heuristically whether or not to add a CNF-encoding of the solution-improving constraint
$\mathcal{O}\leq v^{*}-1.$
For many CNF encodings of pseudo-Boolean constraints, it is well-known how to get certification Vandesande (2023); Gocht et al. (2022); Berg et al. (2024a). One notable exception is the CNF encoding based on binary decision diagrams (BDD), which Bofill et al. (2020) recently generalized to so-called multi-valued decision diagrams (MDD) and is used by MaxCDCL. Most of the interesting questions from a proof logging perspective already show up for BDDs, so to keep the presentation more accessible, we focus on that case and afterwards briefly discuss how this generalizes to MDDs. Formal details and proofs are included in the supplementary material.
@@
The standard way to create a reduced and ordered BDD for
$\sum_{i=1}^{n}v_{i}b_{i}\leq v^{*}-1$
is to first recursively descend and create BDDs $\eta_{\mathbf{t}}$ and $\eta_{\mathbf{f}}$ for
$\sum_{i>1}^{n}v_{i}b_{i}\leq v^{*}-1-a_{1}\text{\quad and}$
$\sum_{i>1}^{n}v_{i}b_{i}\leq v^{*}-1$
respectively, and then combine them into the node $\mathrm{bdd}(x_{1},\eta_{\mathbf{t}},\eta_{\mathbf{f}})$. Using memoization, the BDD is always kept reduced: if a node is being created with the same label, and the same two children as an existing node, the existing node is returned instead; if $\eta_{\mathbf{t}}=\eta_{\mathbf{f}}$, then $\eta_{\mathbf{t}}$ is returned.
@@
$v_{\eta}$ $\Rightarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq l\text{\quad and}$
$v_{\eta}$ $\Leftarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq u.$
@@
$v_{\eta}$ $\Leftrightarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq l\text{\quad and}$
$v_{\eta}^{\prime}$ $\Leftrightarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq u.$
@@ 0.5 Experiments
We implemented proof logging in the MaxCDCL solver as described by Coll et al. (2025b). [^4] On top of the basic algorithm described above, the experiments include proof logging for a range of techniques included in MaxCDCL Coll et al. (2025b), namely clause vivification, equivalent and failed literal detection, clause subsumption and simplification due to unit propagation. Each solver call was assigned a single core on a 2x32-core AMD EPYC 9384X. The time and memory limits are enforced by runlim. [^5] https://github.com/arminbiere/runlim As benchmarks, we used the instances of the MaxSAT Evaluation 2024 Berg et al. (2024b). The implementation, together with the raw data from the experiments can be found on Zenodo Vandesande et al. (2025).
@@
A scatter plot evaluating the overhead in the solving time due to proof logging is shown in Figure 1. The experiments ran with a time limit of 1 hour and a memory limit of 32GB. Out of the $701$ instances that were solved without proof logging, six instances could not be solved with proof logging. Considering instances solved both with and without proof logging, the median overhead of proof logging is $19\%$, which shows that for most instances, proof logging overhead is manageable. However, we see that for $10\%$ of the instances, MaxCDCL with proof logging takes more than $4.61$ times the time to solve the instance using MaxCDCL without proof logging. This is in line with earlier work on certifying PB-to-CNF encodings using VeriPB, such as the work of Berg et al. (2024a). Preliminary tests indicate as possible explanation for this phenomenon that writing the defining constraints for a node in an MDD is linear in the number of objective literals. In practice, this leads to large overhead for instances with a large number of objective literals. Investigating possible ways to circumvent this is part of planned future work.
@@
While optimizing the checking time for the produced proofs was out of scope for this work, we also evaluated it by running the VeriPB proof checker VeriPB with a time limit of 10 hours and a memory limit of 64GB. Figure 2 shows a scatter plot comparing the solving time with proof logging with the proof checking time. Out of $695$ produced proofs, $485$ were checked successfully; for the other $210$ instances, the proof checker ran out of time for $202$ instances and out of memory for $8$ instances. In the median proof verification took $42.94$ times the solving time. We believe this checking overhead to be too high for making this usable in practice. However, there are several ways in which this can be improved. On the one hand, a new proof checker is currently under development for improved performance PBOxide (2025). Simultaneously, there are ongoing efforts for building a proof trimmer for VeriPB proofs. The proofs generated by our tools can serve as benchmarks to further guide this ecosystem of VeriPB tools. On the other hand, once the proof checker is in a stable state, we can investigate which potential optimizations to the generated proofs benefit the proof checking time. One thing that comes to mind is using RUP with hints instead of plain RUP, so that the checker does not need to perform unit propagation. [^6] For more information on these rules, see the VeriPB repository VeriPB .
Figure 2: Comparison of solving time (with proof logging enabled) with proof checking time.
@@ 0.6 Conclusion
As future work, we plan to add proof logging to literal unlocking, a recent technique Li et al. (2025) that improves lower bounds during look-ahead.